# Enhanced DeepONet for 1-D consolidation operator learning: an architectural investigation

### v1

@@ 6 Results
Figure 7 shows the comparative performance of the four DeepONet architectures for a representative case in the test data with a non-uniform $u_{0}$ and a $C_{v}=0.5\ m^{2}/year$. A non-uniform $u_{0}$ is selected as it imposes fluctuating responses early in the consolidation process and enables a more rigorous assessment of the implemented architectures. Moreover, non-uniform $u_{0}$ profiles appear in geotechnical practice (Lovisa and Sivakugan, 2015), for instance, due to staged construction, uneven surface loading, or localized drainage.
@@
Figure 8: Excess pore water pressure (PWP) dissipation evaluated by the different DeepONet architectures on test data at $C_{v}=0.7$ with constant initial excess PWP over the depth. The figures in the first row show the initial excess PWP and the true solution. The subsequent rows show the prediction from each model with the absolute error compared to the true solution, and the plots comparing the true and predicted excess PWP distributions over depth at selected times.
@@
Figure 9: Mean and standard deviation of MSE for each DeepONet architecture across training, validation, and test datasets.
@@
Figure 10: Worst case predictions for Model 3 and Model 4. Note that the figures do not share the same color bar scale for the prediction and error fields.
@@
While the absolute time differences are marginal (order of milliseconds to sub-seconds) for 1-D consolidation (table 1), the results emphasize DeepONet’s potential as a computationally efficient surrogate model. This efficiency could be particularly significant when modeling large-scale or high-dimensional problems where conventional solvers may be computationally costly. Once trained, DeepONet enables rapid inference at any queried spatiotemporal regions without time-stepping across varying initial and material conditions and without retraining.
Table 1: Comparison prediction/execution time (in seconds) for DeepONet architectures and classical numerical solvers.
@@
7 Discussion: Performance of evaluated models and assessment of out-of-distribution data
@@
A key finding in the result is the substantially improved prediction accuracy that Model 3 achieves, which takes the coefficient of consolidation, $C_{v}$, as an input to the trunk net rather than the branch net (i.e., models 1 and 2), which is the classical approach. This improved accuracy is supported by model 3’s architectural similarity to the analytical solution form of the consolidation equation (Equation 5). In Equation 5, $C_{v}$ is an intrinsic component of the dimensionless time factor, $T_{v}=\frac{C_{v}t}{H_{dr}^{2}}$, which directly modulates the exponential decay terms (e.g., $e^{-M^{2}T_{v}}$). These exponential terms essentially act as the time-dependent basis functions in the solution. The trunk net in DeepONet is primarily responsible for learning and representing these basis functions of the solution operator. By providing $C_{v}$ directly as an input to the trunk net alongside the spatiotemporal coordinates $(z,t)$, the architecture better conditions the network to learn these $C_{v}$-dependent basis functions. This architectural alignment with the analytical solution allows Model 3 to more effectively capture the influence of $C_{v}$ on the spataio-temporal evolution of the excess PWP field. From a geotechnical perspective, $C_{v}$ encapsulates the combined effects of soil permeability and compressibility, which are fundamental in predicting the overall consolidation timeline. An architecture that directly learns this influence on the temporal basis functions, like Model 3 does, is therefore more likely to align with the physical principles of the consolidation process. Moreover, Fourier feature embedding effectively improves the accuracy of model 3 (i.e., model 4) (Figure 10), particularly in the early stages of consolidation. By mapping the low-dimensional inputs $(z,t,C_{v})$ into a higher-dimensional space via sinusoidal functions, Fourier feature embedding enriches the input representation and enables the network to more readily approximate functions with high-frequency content or sharp spatio-temporal patterns. This architectural refinement accounts for Model 4’s superior accuracy, especially in challenging scenarios with steep excess PWP gradients. These observations can guide the implementation of future DeepONet architectures in geotechnical problems with similar PDE-driven systems governed by system coefficients with steep functional variations.
@@
We also evaluate the performance of the implemented DeepONet architecture on out-of-training distribution data. First, we assess the solution accuracy of Model 4 for values of $C_{v}$ outside the training range (Figure 12). For this, we create an extra test data set with $C_{v}$ randomly sampled from 0.1 to 1.5 $m^{2}/year$. Within the training range of $C_{v}$ (0.3 to 1.0 $m^{2}/year$, indicated by the shaded region), the model maintains low MSE, typically below $10^{3}\ Pa^{2}$. However, as $C_{v}$ goes into the out-of-distribution region, the MSE shows an increase. This behavior is not unexpected (e.g., (Lu et al., 2021; Ahmed et al., 2024)) because DeepONet is an effective parameter-aware interpolator rather than an extrapolator. This implies that the successful deployment of a DeepONet-based surrogate model requires a careful selection of PDE parameters in the training dataset, which, in turn, requires relevant domain knowledge.
@@
We further evaluate the solution accuracy of Model 4 on function spaces characterized by correlation lengths ranging from 0.2 to 0.8, extending beyond the training correlation length of 0.5. For each correlation length, we construct an additional test dataset comprising 10 realizations of Gaussian random fields as input functions. The lower panel in Figure 13 shows examples of these functions at the correlation lengths 0.2, 0.5, and 0.8. Consistent with prior work on DeepONet extrapolation (Lu et al., 2021), the prediction error increases when testing on functions with correlation lengths smaller than the 0.5 used in training. However, the prediction remains within a reasonable range; the mean MSE at a correlation length of 0.2 (approximately $6\times 10^{3}\ Pa^{2}$) is still lower than what we discussed in Figure 10(a), which exhibits an accurate solution except for the early times. Notably, the error does not increase for correlation lengths greater than 0.5, but decreases. This stability likely stems from the fact that GRFs with larger correlation lengths are smoother and more similar to the training function space, which spans 0.5 GRFs and uniform excess PWP functions. As a result, the model effectively interpolates between these GRFs and the constant functions. Along with Figure 12, the result in Figure 13 highlights the importance of careful DeepONet training data design based on domain knowledge.

### v2

@@ 6 Results
Figure 7 shows the comparative performance of the four DeepONet architectures for a representative case in the test data with a non-uniform $u_{0}$ and a $C_{v}=0.5\ m^{2}/year$. A non-uniform $u_{0}$ is selected as it imposes fluctuating responses early in the consolidation process and enables a more rigorous assessment of the implemented architectures. Moreover, non-uniform $u_{0}$ profiles appear in geotechnical practice (lovisa2015non_uniform_u0), for instance, due to staged construction, uneven surface loading, or localized drainage.
@@
Figure 8: Excess pore water pressure (PWP) dissipation evaluated by the different DeepONet architectures on test data at $C_{v}=0.7\ m^{2}/year$ with constant initial excess PWP over the depth. The figures in the first row show the initial excess PWP and the true solution. The subsequent rows show the prediction from each model with the absolute error compared to the true solution, and the plots comparing the true and predicted excess PWP distributions over depth at selected times.
@@
Figure 9: Mean and standard deviation of MSE for each DeepONet architecture across training, validation, and test datasets in physical scale.
@@
Figure 10: Worst case predictions for Model 3 and Model 4. The plots share the y-axis. The figures do not share the same color bar scale for the prediction and error fields.
@@
Oscillatory patterns in the loss histories Figure 11 are often observed in DeepONet and other deep learning models (e.g., supplementary materials from lu2021learning and wang2021learning_pideeponet). These oscillations can be attributed to the stochastic nature of mini-batch gradient updates in the AdamW optimizer, particularly when using relatively small batch sizes and learning rates. Such fluctuations reflect stochastic variations preventing convergence to non-optimal local minima (lu2023benign_lr_large). In general, adjusting the batch size or learning rate, or applying a learning-rate decay schedule, could help smooth the training curves (smith2017don_decay_lr). In this study, training was stopped at approximately 800 epochs due to GPU runtime constraints on the computing cluster, and the model at this epoch was used for consistent comparisons across architectures. While no explicit early-stopping criterion was used, we monitored training and validation losses throughout training to assess convergence. Convergence was evaluated from the validation-loss trajectory. Across all models, improvements diminish in later epochs, and we observe no systematic separation between training and validation losses or evidence of significant loss instability.
The training time for models 1 to 4 is approximately 32, 44, 31, 35 minutes using NVIDIA RTX 4090 GPU. The elevated training time for Model 2 reflects the additional network introduced for the merge net. The additional Fourier feature embedding operation in Model 4 did not increase the training time significantly.
@@
While the absolute time differences are marginal (order of milliseconds to sub-seconds) for 1D consolidation (table 1), the results emphasize DeepONet’s potential as a computationally efficient surrogate model. This efficiency could be particularly significant when modeling large-scale or high-dimensional problems where conventional solvers may be computationally costly. Once trained, DeepONet enables rapid inference at any queried spatiotemporal regions without time-stepping across varying initial and material conditions and without retraining.
Table 1: Comparison prediction/execution time (in seconds) for DeepONet architectures and classical numerical solvers for 1D case.
@@
6.1 Application to 3D Consolidation
We further extend the application of Model 4 to 3D consolidation scenarios; the motivation is to have a scalability assessment. The governing equation for 3D consolidation is:
$\frac{\partial u}{\partial t}=C_{v}\left(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}+\frac{\partial^{2}u}{\partial z^{2}}\right)$ (12)
Here, $u$ is defined in a 3D space with $u(x,y,z,t)$ at time $t$. In this case, fully drained boundary conditions ($u=0$) are applied on all faces of a cubic domain.
The numerical formulation for the 3D case follows the same finite difference approach used in the 1D consolidation problem. The domain is uniformly discretized into $N_{x}$, $N_{y}$, and $N_{z}$ nodes along the $x$, $y$, and $z$ directions, respectively, with grid spacings $\Delta x$, $\Delta y$, and $\Delta z$. Applying a second-order central finite-difference approximation in each direction, Equation 12 can be expressed for an internal node $(i,j,k)$ as:
$\frac{du_{i,j,k}}{dt}=C_{v}\left(\frac{u_{i+1,j,k}-2u_{i,j,k}+u_{i-1,j,k}}{\Delta x^{2}}+\frac{u_{i,j+1,k}-2u_{i,j,k}+u_{i,j-1,k}}{\Delta y^{2}}+\frac{u_{i,j,k+1}-2u_{i,j,k}+u_{i,j,k-1}}{\Delta z^{2}}\right)$ (13)
Similar to the 1D case, we use RK methods to solve the discretized ODE system using the solve_ivp function from the SciPy library
We generate N = 20,000 and N = 2,000 functions and coefficients for training and validation, respectively, and test the trained model on 100 unseen functions and coefficients, with no overlap between the three sets. Specifically, the initial excess pore water pressure, $u(x,y,z,0)$, varies spatially in the $x$ and $y$ directions and is modeled using GRF, as in the 1D case. The GRF is defined with $\sigma_{r}^{2}=1\ \text{kPa}^{2}$ and $l=0.30$ in $x$ and $y$ directions. $\mu_{r}$ is uniformly sampled from 10 to 20 kPa. The input function is represented by $m=50\times 50$ equally spaced locations distributed over the $x$-$y$ plane (i.e., $\{x_{\ell}\}_{\ell}^{m=50\times 50}$). The corresponding solutions are sampled at P=10,000 locations uniformly, (i.e., $\{y_{j}^{(i)}\}_{j=1}^{10,000}$), in the four-dimensional space $(x,y,z,t)$. As a result, we create $20,000\times 10,000$ training data. $C_{v}$ is sampled uniformly from the range $0.3$-$1.0\ m^{2}/year$. The validation set is constructed in the same manner, with P=10,000 output evaluations per input function, resulting in $2,000\times 10,000$ pairs. The spatial domain spans the normalized $x$, $y$, $z$ and $t$, i.e., $\frac{x,\ y,\ z}{H_{dr}}\in[0,2]$ and normalized time $T_{v}\in[0,1]$.
The branch and trunk net uses MLPs with 5 layers and 128 neurons per layer. Their output dimensions for the dot product are set to 128. The Fourier feature embedding uses $m_{f}=50$ and $\sigma_{f}=1.0$. We use the AdamW optimizer, with a learning rate of 1.0e-4, a batch size of 16,834. Compared to the 1D case in Model 4, this setting uses a larger MLP and a larger batch size during training to achieve higher data throughput. We train the model for 70 epochs, taking about 11 hours using the same GPU used in the 1D case.
We evaluate the learned solution operator $\mathcal{G}_{\Theta}(u^{(i)})(y)$ at a $50^{4}$ uniformly spaced grid of $y(x,y,z,t)$ points across the output domain. Figure 12 shows the result with $C_{v}=0.50\ m^{2}/year$ at $T_{v}=0,\ 0.10,\ 0.40$ (results for other $C_{v}$ are similar). The first row shows the input function $u_{0}$. DeepONet captures the overall spatio-temporal PWP dissipation response successfully with an average MSE of 1.07e+03 $Pa^{2}$. The prediction shows the largest overall error at the very first timestep with the maximum absolute error of 4.13e+2 $Pa$. Compared to the 1D case in Figure 7, this value is slightly larger than Model 4’s maximum absolute error (2.97e+2 $Pa$) and smaller than the maximum absolute error from Model 3 (8.69e+2 $Pa$). For the later times, the prediction errors become almost negligible.
Figure 12: Excess PWP dissipation evaluated by the 3D DeepONet architectures on test data at $C_{v}=0.5\ m^{2}/year$ with spatially varying initial excess PWP under full drainage condition. The subsequent rows show the solution field from the finite difference method, prediction from DeepONet, and the absolute error, at the specified $T_{v}$
DeepONets are not restricted to the specific boundary conditions and can be expanded to mixed boundary conditions depending on training data coverage. We prepare an additional consolidation dataset with a bottom impermeable condition with the same data size as the full drainage case to train the model. Figure 13 shows an evaluation result with $C_{v}=0.35\ m^{2}/year$ at $T_{v}=0,\ 0.07,\ 0.28$. Similar to the case with full drainage conditions, the prediction shows a good agreement with the finite difference solver. Although we do not address the other boundary conditions, in principle, they can be accommodated by incorporating boundary-condition indexing variables or additional input channels into the operator formulation.
Figure 13: Excess PWP dissipation evaluated by the 3D DeepONet architectures on test data at $C_{v}=0.35\ m^{2}/year$ with spatially varying initial excess PWP under bottom impermeable condition. The subsequent rows show the solution field from the finite difference method, prediction from DeepONet, and the absolute error, at the specified $T_{v}$
Table 2 shows the mean and standard deviation of MSE for evaluating the full solution field in physical scale across training, validation, and test datasets for both full drainage and bottom impermeable boundary conditions. We can see that these MSEs span between the values of Model 3 and Model 4 in the 1D case (see Figure 9), with significantly smaller values than Model 1 and 2. Given the greater complexity and larger sampling space of the 3D problem compared to the 1D case, Model 4 demonstrates good performance across the considered 3D scenarios.
Table 2: Model performance under different drainage conditions.
Dataset Full Drainage Bottom Impermeable
Mean MSE ($\mathrm{Pa}^{2}$) Std (Pa) Mean MSE ($\mathrm{Pa}^{2}$) Std (Pa)
Training $7.65\times 10^{2}$ $3.01\times 10^{1}$ $5.49\times 10^{2}$ $2.45\times 10^{1}$
Validation $1.09\times 10^{3}$ $2.90\times 10^{1}$ $5.49\times 10^{2}$ $2.62\times 10^{1}$
Test $7.78\times 10^{2}$ $3.51\times 10^{2}$ $5.13\times 10^{2}$ $2.61\times 10^{2}$
Table 3 compares the computation time required to generate a full 3D solution field between the trained DeepONet and the explicit numerical solver (RK45). DeepONet predicts the complete 3D field almost instantaneously (about 0.1 s), whereas the classical solver requires over 120 s on average, achieving a speedup of more than three orders of magnitude. The implicit BDF solver was unable to produce results, as the computer memory ran out due to the prohibitively large Jacobian matrix in the 3D configuration.
Table 3: Comparison of prediction/execution time (in seconds) for DeepONet and classical numerical solvers for the 3D case.
Method Mean (s) Std (s)
Model 4 (Standard) 0.101 0.019
Explicit solver (RK45) 126.63 4.30
Figure 14 further illustrates the performance of the trained DeepONet in 3D scenarios by showing the evolution of the spatially averaged degree of consolidation $U(t)$ under uncertain conditions, comparing DeepONet predictions with those from the finite difference solver (reference solution). Figure 14 also highlights a potential practical application of DeepONet because $U(t)$ is a key engineering parameter to make decisions, and quantifying its uncertainty is of interest. In engineering applications, pore pressure instrumentation often reveals spatially variable piezometric regimes (tang2025consol_non_uniform_loads; darrag1993consolidation_stochastic_initial), which can be approximated by random fields using geostatistical techniques (e.g., macedo2024cptu), making them well suited for DeepONet-based surrogate modeling. Moreover, engineering assessments commonly use mean values of $C_{v}$, varied over lower and upper bounds to reflect parameter uncertainty (macedo2024cptu). We generate 1,000 realizations of the initial excess PWP using GRF with $\sigma_{r}=1\ kPa^{2}$, $\mu_{r}=15\ kPa$, and $l=0.3$. $C_{v}$ is sampled from a normal distribution with mean $0.5~\mathrm{m^{2}/year}$ and a coefficient of variation of 0.1. For each realization, the corresponding $U(t)$ is computed by using the dissipated pore pressures at a given time and averaging over the calculation domain; this process is described in classical geotechnical books (e.g., holtz1981introduction_geotech). The $U(t)$ variations at a given time are used to approximate its statistical distribution and plot the uncertainty range in Figure 14. DeepONet accurately reproduces uncertainty propagation, yielding $\pm 2\sigma$ bounds nearly identical to those from the finite difference solver while retaining the computational efficiency reported in Table 3. For example, at $t=0.1$ years, the mean and standard deviation of $U$ from the numerical solver are 0.8733 and 0.0227, respectively, whereas DeepONet yields 0.8731 and 0.0228, demonstrating an excellent quantitative agreement. These results indicate that DeepONet can efficiently and reliably quantify uncertainty under variable input conditions, offering a promising surrogate modeling strategy for complex consolidation problems. In practice, DeepONet could be integrated with modern sensing technologies (e.g., anjana2025fiber_pwp and hottges2023fiber_novel) that collect dense pore-pressure measurements, together with robust subsurface characterization methods to estimate bounds on $C_{v}$. Such a framework has the potential to enable computationally efficient, large-scale, near real-time consolidation monitoring, where DeepONet’s speed offers substantial advantages over traditional solvers.
Figure 14: Stochastic evolution of the degree of consolidation $U(t)$ obtained from the finite difference solver (left) and DeepONet (right), showing close agreement in mean trends and $\pm 2\sigma$ uncertainty bounds.
7 Discussion
7.1 Analytically guided architecture performance
A key finding is the substantially improved prediction accuracy achieved by Model 3, in which the coefficient of consolidation $C_{v}$ is provided to the trunk net rather than the branch net (as in Models 1 and 2). This improvement is supported by the close correspondence between Model 3 and the analytical solution structure of the consolidation equation (Equation 5), given by
$u(z,t;C_{v})=\sum_{a=0}^{\infty}\underbrace{\frac{2u_{0}}{A}}_{b_{a}(u_{0})}\;\underbrace{\sin\!\left(\frac{Az}{H_{\mathrm{dr}}}\right)\exp\!\left(-A^{2}\frac{C_{v}t}{H_{\mathrm{dr}}^{2}}\right)}_{\phi_{a}(z,t,C_{v})}\;\approx\;\sum_{k=1}^{q}b_{k}(u_{0})\,t_{k}(z,t,C_{v})$ (14)
Where $b_{a}(u_{0})$ depends only on the initial condition, and $\phi_{a}(z,t,C_{v})$ are spatio-temporal basis functions parametrized by $C_{v}$. The right-hand side represents the corresponding DeepONet approximation, in which $b_{k}$ and $t_{k}$ are learned by the branch and trunk networks, respectively. Notably, $C_{v}$ appears only in the exponential decay term through the dimensionless time factor $T_{v}=\frac{C_{v}t}{H_{\mathrm{dr}}^{2}}$, directly controlling the temporal evolution of the basis functions. Embedding $C_{v}$ in the trunk net therefore enables the learned basis functions $t_{k}(z,t,C_{v})$ to adapt to different consolidation rates in a manner consistent with the analytical solution. This architectural alignment allows Model 3 to more effectively capture the influence of $C_{v}$ on the spatio-temporal evolution of the excess PWP field.
@@ 7.1 Analytically guided architecture performance
From a geotechnical perspective, $C_{v}$ encapsulates the combined effects of soil permeability and compressibility, which are fundamental in predicting the overall consolidation timeline. An architecture that directly learns this influence on the temporal basis functions, like Model 3 does, is therefore more likely to align with the physical principles of the consolidation process. Moreover, Fourier feature embedding effectively improves the accuracy of Model 3 (i.e., Model 4) (Figure 10), particularly in the early stages of consolidation. By mapping the low-dimensional inputs $(z,t,C_{v})$ into a higher-dimensional space via sinusoidal functions, Fourier feature embedding enriches the input representation and enables the network to more readily approximate functions with high-frequency content or sharp spatio-temporal patterns. This architectural refinement accounts for Model 4’s superior accuracy, especially in challenging scenarios with steep excess PWP gradients. These observations can guide the implementation of future DeepONet architectures in geotechnical problems with similar PDE-driven systems governed by system coefficients with steep functional variations.
7.2 Out-of-distribution response
@@
We also evaluate the performance of the implemented DeepONet architecture on out-of-training distribution data focusing on 1D case. First, we assess the solution accuracy of Model 4 for values of $C_{v}$ outside the training range (Figure 15). For this, we create an extra test data set with $C_{v}$ randomly sampled from 0.1 to 1.5 $m^{2}/year$. Within the training range of $C_{v}$ (0.3 to 1.0 $m^{2}/year$, indicated by the shaded region), the model maintains low MSE, typically below $10^{3}\ Pa^{2}$. However, as $C_{v}$ goes into the out-of-distribution region, the MSE shows an increase. This behavior is not unexpected (e.g., (lu2021learning; ahmed2024deeponet_structure)) because DeepONet is an effective parameter-aware interpolator rather than an extrapolator. This implies that the successful deployment of a DeepONet-based surrogate model requires a careful selection of PDE parameters in the training dataset, which, in turn, requires relevant domain knowledge.
@@
We further evaluate the solution accuracy of Model 4 on function spaces characterized by correlation lengths ranging from 0.2 to 0.8, extending beyond the training correlation length of 0.5. For each correlation length, we construct an additional test dataset comprising 10 realizations of Gaussian random fields as input functions. The lower panel in Figure 16 shows examples of these functions at the correlation lengths 0.2, 0.5, and 0.8. Consistent with prior work on DeepONet extrapolation (lu2021learning), the prediction error increases when testing on functions with correlation lengths smaller than the 0.5 used in training. However, the prediction remains within a reasonable range; the mean MSE at a correlation length of 0.2 (approximately $6\times 10^{3}\ Pa^{2}$) is still lower than what we discussed in Figure 10(a), which exhibits an accurate solution except for the early times. Notably, the error does not increase for correlation lengths greater than 0.5, but decreases. This stability likely stems from the fact that GRFs with larger correlation lengths are smoother and more similar to the training function space, which spans 0.5 GRFs and uniform excess PWP functions. As a result, the model effectively interpolates between these GRFs and the constant functions. Along with Figure 15, the result in Figure 16 highlights the importance of careful DeepONet training data design based on domain knowledge.
@@
While this study primarily focuses on evaluating DeepONet’s potential as an accurate solver for consolidation problems, through an architectural exploration, future work can address extrapolation limitations by incorporating physics-informed and adaptive training strategies, i.e., including domain expertise to ensure plausible ranges of $C_{v}$ and spatial pore pressure distributions are represented into the trained DeepONet for different geologies of interest.
Alternatively, a physics-informed DeepONet (wang2021learning_pideeponet), in which the governing consolidation equation (Equation 4) is incorporated as a soft constraint in the loss function, could also enhance physical consistency and improve generalization outside the training range. In this setting, automatic differentiation is used to compute the residual of the PDE at collocation points in the spatio-temporal domain, and an additional physics-based loss term can be introduced to penalize violations of the PDE and boundary conditions, similar to the approaches from PINNs (raissi2019pinn). For the present consolidation problem, such a formulation could explicitly enforce diffusion-driven dissipation and pore pressure decay with their parameter dependence, potentially improving out-of-distribution errors. However, balancing data loss and physics-based constraints remains challenging, and improper weighting of them may lead to biased solutions or slow convergence (faroughi2022sciml).
In addition, active learning frameworks offer a promising strategy for adaptively enriching the training dataset in regions associated with high uncertainty or large prediction errors (guan2022adaptive_active; tharwat2023survey_active). For consolidation problems, this could be implemented by iteratively identifying ranges of $C_{v}$, correlation lengths, or early-time intervals where models exhibit elevated MSE, and selectively generating additional numerical simulations in these regimes. The newly sampled cases can then be incorporated into subsequent training rounds to improve coverage of challenging regions in the input and parameter spaces. Such an approach is expected to enhance generalization while minimizing the computational cost of data generation without full retraining. Nevertheless, active learning presents complexities in uncertainty estimation and sample selection, as its effectiveness is often linked to the reliability of the metrics used to gauge model confidence (winovich2025active).
7.3 Performance in the context of other operator-learning neural networks
Besides DeepONet, there are also other operator-learning neural networks such as Fourier neural operator (FNO) (li2020fourierNOs), and transformer-based neural operators (TBNOs) (hao2023transformer_no) that could be potentially considered for the consolidation problem. FNO learns operators using spectral convolutions: it maps input functions to Fourier space, applies learnable filters to selected frequency modes, and transforms back to the physical domain, enabling efficient learning of smooth global dependencies of the domain. A detailed description is provided in li2020fourierNOs. TBNOs extend the attention mechanisms (vaswani2017attention) of transformers, which were originally developed to learn sequential dependencies, to operator learning in space and time. hao2023transformer_no provides additional details. Despite their increased learning capacity, TBNOs typically require more computational resources due to their use of complex neural network structures such as transformers (hao2023transformer_no; li2022transformer_no). In this study, we limit the additional exploration of alternative operator-learning neural networks to FNOs, which provide a well-established and computationally practical baseline comparable to DeepONet.
We construct the FNO on a $100\times 100$ uniform grid spanning the normalized depth and time, $\frac{z}{H_{dr}}\in[0,1]$ and $T_{v}=\frac{C_{v}t}{H_{dr}^{2}}\in[0,2]$, the same range as DeepONet. Each grid point is represented by four input channels $C_{v}$, $z$, $t$, and $u_{0}$. The architecture comprises four spectral convolution layers with a hidden channel dimension of 8, followed by an output layer that produces $100\times 100$ field representing the excess PWP $u(z,t)$. Thus, a single training sample consists of an input tensor of shape $100\times 100\times 4$ and a corresponding target solution tensor of shape $100\times 100\times 1$.
The model is trained with the AdamW optimizer using a learning rate of 0.01, a batch size of 256, and 800 epochs. The training dataset contains 40,000 samples, and both the validation and test sets contain 5,000 samples. These training hyperparameters align with those used for Model 4. The resulting MSEs on the training, validation, and test sets are 116.49 $\text{Pa}^{2}$, 119.21 $\text{Pa}^{2}$, and 481.69 $\text{Pa}^{2}$ (with standard deviation of 153.63 $\text{Pa}^{2}$, 155.82 $\text{Pa}^{2}$, and 186.15 $\text{Pa}^{2}$), respectively. Figure 17 presents the worst-case prediction, like we showed in Figure 10.
Figure 17: Worst case predictions from FNO with $C_{v}=0.3009\ m^{2}/year$. The plots share the y-axis.
Although the FNO achieves lower training and validation MSEs than Model 4 (the enhanced DeepONet), its test MSE (481.69 $\text{Pa}^{2}$) does not show a meaningful difference from Model 4’s test MSE (508.96 $\text{Pa}^{2}$ Figure 9). This indicates that, in our problem setting, while the FNO fits the training distribution more effectively, its improved representation performance does not translate into substantially better generalization to unseen functions.
A contributing factor is the effective volume of supervised data each model receives. While both architectures are trained with 40,000 input functions, the supervision density differs due to their inherent architectural design: DeepONet observes only $10\times 10$ sparse solution samples per function, whereas FNO is trained on dense $100\times 100$ regular solution fields spanning over the solution domain. Thus, FNO effectively receives far more labeled output data per function, which partly explains its lower prediction errors.
The training efforts also vary. Training FNO took approximately 165 minutes, over four times longer than Model 4 (around 40 minutes). This difference shows that the DeepONet can achieve comparable generalization accuracy with lower training costs for the consolidation problem being examined. Sensitivity to architecture and problem type might be a factor, so this is not a general statement for these operator learning frameworks. Such comparisons are beyond the scope of this study and can be conducted for a broader set of geotechnical problems in future efforts.
DeepONet’s inference time and memory usage scale with the number of query points, whereas FNO’s computational cost primarily depends on the output grid resolution due to their inherent architectural differences. In our test setting, with a $50\times 50$ output field, both DeepONet (Model 4) and FNO enable fast evaluation once trained, with inference costs that are negligible compared to classical numerical solvers. Specifically, Model 4 requires approximately $2.43\,\mathrm{s}$ to generate predictions for the entire test set (i.e., 500 solution fields), while FNO requires $4.12\,\mathrm{s}$. In terms of memory consumption, DeepONet uses about $120\,\mathrm{MB}$ of GPU memory, whereas FNO requires approximately $374\,\mathrm{MB}$. This difference is likely due to the relatively small size of Model 4, which allows it to outperform the global Fourier transforms employed by FNO on a $50\times 50$ grid.
Even though DeepONet and FNO performance were comparable for the consolidation problem, there are practical advantages for DeepONet over FNO. FNO assumes data defined on uniform Cartesian grids in order to leverage Fast Fourier Transform (FFT)-based convolutions. Extensions to irregular geometries or non-uniform meshes typically require additional preprocessing, mapping, or architectural modifications (li2023fourier_geom). DeepONet, by contrast, does not impose grid regularity constraints and can naturally accommodate non-uniform meshes or irregular output sampling through its pointwise trunk inputs at queried locations. This allows DeepONet to be more flexible in irregular geometries (lu2021learning; faroughi2022sciml). Additionally, DeepONet’s flexible branch-trunk architecture accommodates diverse neural modules (e.g., MLPs, convolution neural networks (mei2024cnn_don), graph networks (zhang2025graph_don), and attention mechanisms (chen2025attension_don)), facilitating hybrid designs and multimodal operator inputs. The training for DeepONet was also shorter for the configuration in this study. Finally, DeepONet is also supported by the theoretical foundation, universal approximation theorem for nonlinear operators in infinite-dimensional spaces (chen1995universal; lu2021learning).
7.4 DeepOnet in the context of classical surrogate modeling
Classical surrogate modeling approaches, such as projection-based reduced-order models (ROMs) and polynomial chaos expansions (PCEs), are commonly used to approximate high-fidelity simulations, particularly for problems with low-dimensional parametric inputs or limited solution variability. However, their accuracy and computational efficiency often deteriorate as the dimensionality of the parameter space and the spatial complexity of the solution increase (benner2015survey-rom). Moreover, projection-based ROMs rely on a fixed spatial discretization, which limits their flexibility when transferring models across varying meshes during inference. In contrast, DeepONet is not restricted to a prescribed discretization, enabling greater adaptability to heterogeneous spatial representations (dummer2026ronomreducedorderneuraloperator).
To illustrate the performance of DeepONet in the context of classical ROMs, we compare DeepONet (Model 4) with a proper orthogonal decomposition (POD)-based ROM (xiao2017parameterized; gao2024machine-rom) in a three-dimensional setting. The ROM is constructed by computing a truncated POD basis via singular value decomposition of full-field solution snapshots. Retaining 50 modes yields a set of orthonormal spatial basis functions, $\boldsymbol{\Phi}\in\mathbb{R}^{n\times 50}$. An MLP with 256 hidden units maps the flattened initial excess pore water pressure $u_{0}$, the coefficient of consolidation $C_{v}$, and time $t$ to the corresponding POD coefficients $\boldsymbol{c}\in\mathbb{R}^{50}$. In this ROM formulation, the spatial basis $\boldsymbol{\Phi}$ is precomputed during training and remains fixed during inference, while only the coefficient mapping is learned. By contrast, DeepONet simultaneously learns both the basis functions (through the trunk network) and the associated coefficients (through the branch network), with the trunk network evaluated at each output location $(x,y,z,t)$. This operator-learning formulation enables DeepONet to represent more flexible and adaptive solution manifolds.
We evaluate the ROM on 10 uniformly sampled test cases under the same conditions used for DeepONet. The ROM attains a mean field mean-squared error (MSE) of $1.528\times 10^{6},\mathrm{Pa}^{2}$, with a mean prediction time of 0.105 s per full three-dimensional field. A representative prediction is shown in Figure 18. While the computational cost is comparable to that of DeepONet (Table 3), the ROM consistently exhibits larger prediction errors (Figure 12), particularly at early times and in regions where the initial excess pore water pressure exhibits strong spatial gradients.
Several caveats apply to this comparison. Data equivalence is defined in terms of total storage requirements: the ROM is trained using 250 full-field input–output pairs, which occupy approximately the same storage as the DeepONet training dataset. However, the two approaches rely on fundamentally different data representations. The ROM requires dense full-field snapshots for offline basis construction, whereas DeepONet is trained using sparsely sampled pointwise evaluations, complicating a strictly equitable data comparison. Furthermore, the effective model capacities are governed by distinct and non-commensurable hyperparameters, including the number of retained POD modes and the MLP architecture for the ROM, versus the branch-trunk network widths and Fourier feature settings for DeepONet. A systematic exploration of these factors is beyond the scope of this study and represents an important direction for future work.
Figure 18: Excess PWP dissipation evaluated by the ROM on test data at $C_{v}=0.35\ m^{2}/year$ with spatially varying initial excess PWP. The subsequent rows show the solution field from the finite difference method, prediction from ROM, and the absolute error, at the specified $T_{v}$